Tohban Report 2015-05-27
From RHESSI Wiki
|Start Date:||20 May 2015|
|End Date:||27 May 2015|
|Next Tohban:||Nicole Duncan|
|List all reports|
Solar activity was really low this week. Currently there are four very little active regions on the disk with very low probability to host big flares. Three of them are rotating off of disk in the next couple of days. Solar activity is expected to be low next days.
How many GOES flares occurred?
Flares above B, C, M, X class were 18 10 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 13 8 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 0 0 0
There were RHESSI flares/GOES flares 100 / 28 over the time range 18-May-15 25-May-15
The SSR fill level continued high most of the week but during last couple of days has shown a good decrease. Currently SSR reaches a maximum of around 30% and is emptying at the end of daily passes.
Normal as far as the Tohban can tell.
No data gaps this week
On May 21 starting at 19:23 UTC we made a few changes to detectors 2, 4 and 6. We raised thresholds on the Front Slow on all of them as follow:
|Threshold [G2]||Front Slow|
|Threshold [G4]||Front Slow|
|Threshold [G6]||Front Slow|
In addition, yesterday (May 26) starting at 19:11 UT we lowered the HV on detector 6 by 30 counts as follows:
With this last change detector 6 has the lowest voltage among all the detectors. Despite this, detector 6 has remained still segmented. During that pass we also carried out a couple of increasing of D5 and D6 Rear Fast thresholds.
|Threshold [G5]||Rear Fast|
|Threshold [G6]||Rear Fast|
After all this changes all detectors seems mostly to be stable. Detector 3 is showing a relatively low value on its live-time for its front segment in addition to a light leak. Next Tohban should consider changing its threshold.
- Last Friday May 22 starting at 16:00 UT we stopped taking night-times in order to improve the SSR behaviour. Now that things are quiet down, next Tohban should consider turning on the night events.
Temperatures are lightly higher than a couples of weeks ago. We need to discuss deeply this issue during the meeting.
Cold Tip Temperature 1 109.716 K Cold Tip Temperature 2 107.987 K Cold Plate Temperature 1 132.311 K Cold Plate Temperature 2 130.664 K
Light leak on front sides of detectors 2 and 3.
Now that the SSR fill level is quiet down, next Tohban should consider turning on the night events.
|Night time data (fronts)||No nighttime events|
|Night time data (rears)||No nighttime events|
|Require extra passes?||No|
|Requirement for moving pointer?||No|